OpenStudy (anonymous):

the widths of 2 similar rectangles are 16 cm and 14 cm. whats the ratio of the area?

OpenStudy (anonymous):

b1/b2=16/14=l1/l2. l is the length . so the area1/area2=14^2/16^2 = 49/56

OpenStudy (anonymous):

similar rectangle means their length and width are in same ratio....

OpenStudy (anonymous):

if each dimension of a rectangle family room is 3.5 tmme larger than a rectangular bedrrom. if the cost 312 to carpet the bedroom, how much would it cost to carpet the family room with the same carpet

OpenStudy (anonymous):

let length is l and width is d for bedroom so that for family room is 3.5l and 3.5 d area of the family room 3.5^2ld and for bedroom ld but you have for bedroom the cost is 312 so for familyroom it is (3.15)^2*312 now u can go on.......

OpenStudy (anonymous):

how do u find area for a kite

OpenStudy (anonymous):

is the ans like 16/14 ie 8/7????i thnk so

OpenStudy (anonymous):

the area of kite is given by the half of the product the diagonals. as it is rhombus.

OpenStudy (anonymous):

why are you saying that area would be in ration 8/7....it is given that the rectangle are similar.. i have attached a file for the similar rectangles check it...

OpenStudy (anonymous):

ans is 8/7

OpenStudy (anonymous):

no anns is 64/49

OpenStudy (anonymous):

why are you saying that are you considering that length is same........??????????

OpenStudy (anonymous):

like similar triangles the similar rectangles should exhibit same property of the side ratio.........